OpenStudy (anonymous):

Britney is preparing boxes of assorted chocolates. Chocolate covered peanuts cost $7 per pound. Chocolate covered caramel calls $6.50 per pound. The boxes of assorted candies contain five more pounds of peanut candies than caramel candies. if the total amount sold was $575, and many pounds of each candy were needed to make the boxes?

OpenStudy (anonymous):

45 lbs of peanut candies and 40 lbs of carmel candies

OpenStudy (anonymous):

NO GIVING OUT ANSWERS.

OpenStudy (anonymous):

@justsmile531 you keep beating me.. -.-

OpenStudy (anonymous):

\[P=C+5\]\[7P+6.5C=575\]\[7(c+5)+6.5c=575\]\[13.5c+35=575\]\[c=40\]\[P=40+5\]\[p=45\] 40lbs Caramel, 45lbs Peanut.